OpenStudy (accidentalaichan):

what is the product of 4 sqrt 6 * 8 sqrt 14?

OpenStudy (solomonzelman):

\(\Large\color{darkgreen}{ 4\sqrt{6} \times 8\sqrt{14} }\) \(\Large\color{darkgreen}{ 4 \times \sqrt{6} \times 8 \times \sqrt{14} }\) \(\Large\color{darkgreen}{ 4 \times 8 \times \sqrt{14}\times \sqrt{6} }\) \(\Large\color{darkgreen}{ 4 \times 8 \times \sqrt{14\times 6} }\)

OpenStudy (solomonzelman):

can you finish the problem ?

OpenStudy (accidentalaichan):

32 \[\sqrt{84}\]? Now, simplify?

OpenStudy (solomonzelman):

\(\Large\color{darkgreen}{ 32 \sqrt{84} }\)

OpenStudy (solomonzelman):

You can actually simplify more, \(\Large\color{darkgreen}{ \sqrt{84} =\sqrt{4} \times \sqrt{21} }\)

OpenStudy (accidentalaichan):

Okay, I think it's \[64 \sqrt{42}\]

OpenStudy (solomonzelman):

\(\Large\color{red}{ \bf C }\)\(\Large\color{blue}{ \bf ORRECT ! }\)

OpenStudy (accidentalaichan):

Thanks for your help!

OpenStudy (solomonzelman):

Anytime.

OpenStudy (ipwnbunnies):

Nooo, wait. Don't you mean \[64\sqrt{21}\]?

OpenStudy (ipwnbunnies):

\[32 * \sqrt{4} * \sqrt{21}\] \[32 * 2 * \sqrt{21}\] \[64\sqrt{21}\]